OpenStudy (anonymous):

Solve this quadratic equation using the quadratic formula. x2 + 8x - 5 = 0 i need to show my work!

OpenStudy (jkristia):

try and plug it in to the quadratic formula.

OpenStudy (anonymous):

identify a = b = c =

OpenStudy (anonymous):

ax^2 + bx + c = 0 x^2 + 8x -5 = 0

OpenStudy (anonymous):

\[ax^2+bx+c=0\\ x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

Oh dude I WISH I were still doing this sort of thing!

OpenStudy (anonymous):

I know how to do the quadratic formula I would really just want the answer to compare to mine and see if I got it right.

OpenStudy (jkristia):

so what is your answer?

OpenStudy (anonymous):

a=1 b=8 c=-5 because of the standard form ax^2+bx+c Plug those numbers into a, b, and c for that quadratic formula. Just plug and chug my friend!``

OpenStudy (jkristia):

you can always replace 'x' with your answer and see it you get 0, if you do you got it right.